Please help with the homework:
1.A jar has 7 red marbles,4 black marbles and 1 white marble. Suppose each marble has an equal chance of being picked from the jar.
1.1. find the probability that a black marble is picked?
1.2. what is the probability of picking a red and a white marble?

Answer by math_helper(2461)


1.1. On ONE pick, Pr(black) = 4/(7+4+1) = 4/12 = 1/3
1.2. On TWO picks without replacement,
Pr(Red and White) = (7/12)*(1/11) + (1/12)*(7/11) = 14/132 = 7/66